A cannonball shot at 51 m/s at an angle of 38°. What will be the maximum height that the cannonball reaches? Units?


1
Expert's answer
2021-09-20T14:55:38-0400

The maximum height reached by the cannonball can be found as follows:


"y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(51\ \dfrac{m}{s})^2\cdot sin^238^{\circ}}{2\cdot9.8\ \dfrac{m}{s^2}}=50.3\ m."